OpenStudy (anonymous):

e^(x) - 3e^(-x) - 2 = 0 Having a little trouble solving for x.

OpenStudy (anonymous):

e^(x) - 3e^(-x)=2 Ln(e^(x) - 3e^(-x))=Ln2 Ln(e^(x))-3Ln(e^(-x))=Ln2 Ln(e^x/3e^(-x))=Ln2 Ln(e^2x/3)=Ln2 2xLn(1/3)=Ln2 continue

OpenStudy (anonymous):

I'm getting x = ln(2) / 2ln(1/3) , which is wrong according to my book

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

e^(x) - 3e^(-x)=2 Ln(e^(x) - 3e^(-x))=Ln2 Ln(e^(x))-Ln(e^(-3x))=Ln2 Ln(e^x/e^(-3x))=Ln2 Ln(e^4x)=Ln2 4x=Ln2